OpenStudy (anonymous):

What is the exact value of the expression \[\sqrt[5]{5} + \sqrt[3]{20} -\sqrt[6]{3}\]

OpenStudy (anonymous):

Evaluate each one of them using Linear Approximation (Differential Calculus) I hope you know about it.

OpenStudy (anonymous):

can you explain that to me.

OpenStudy (anonymous):

For fifth root of 5.. Let us consider function y = x^ (1/5)

OpenStudy (anonymous):

ok.

OpenStudy (anonymous):

\[\frac{dy}{dx} = \frac{1}{5x^{4/5}}\]